OpenStudy (anonymous):

What is 512 written as a power of a number? A. 92 B. 44 C. 28 D. 29

OpenStudy (anonymous):

please help

OpenStudy (mathstudent55):

Find the prime factorization of 512.

OpenStudy (anonymous):

i dont know how to?

OpenStudy (mathstudent55):

To do that, divide by prime numbers starting with 2 as many times as possible. Then once it is no longer divisible by 2, move on to 3. Keep on going until you get 1.

OpenStudy (mathstudent55):

I'll show you an example.

OpenStudy (anonymous):

can u give me the answer?

OpenStudy (mathstudent55):

Find the prime factorization of 81. The prime numbers are: 2, 3, 5, 7, 11, 13, 17, 19, ... We start by dividing by the smallest prime that divided evenly into 81. 81 is not divisible by 2, so we move to 3. 81/3 = 27 27/3 = 9 9/3 = 3 3/3 = 1 Since we divided 81 4 times by 3, 81 = 3^4

OpenStudy (mathstudent55):

512 is divisible by 2. Just divide it by 2 and keep dividing it by 2 until you get 1. Then count how many times you divided it by 2. That is the exponent on the 2.

OpenStudy (anonymous):

its ok, i got the answer

OpenStudy (anonymous):

thanks!

OpenStudy (mathstudent55):

2^9 ?

OpenStudy (anonymous):

ya

OpenStudy (mathstudent55):

Correct!

OpenStudy (anonymous):

bye!